OpenStudy (anonymous):

if h=85 cm and f=36 cm,what is the lenght of g?

OpenStudy (anonymous):

https://app37.studyisland.com/pics/192060pic9.png here is the picture of the triangle

OpenStudy (anonymous):

a^2+b^2 =c^2 36+g^2 = 85

OpenStudy (anonymous):

solve for g

OpenStudy (anonymous):

would be 49?

OpenStudy (anonymous):